# Robust Instance Optimal Phase-Only Compressed Sensing

### v2

@@ 5 Conclusion
Beyond Theorem 3.3 known from [27], we provided a new set of robustness results. First, dense noise bounded by small enough $\tau_{0}$ (either before or after taking the phases) increments the estimation error by $O(\tau_{0})$, and no algorithm can do substantially better than this. Second, an adversarial $\zeta_{0}$-fraction of sparse corruption increments the error by $\tilde{O}(\sqrt{\zeta_{0}})$. We conjectured that this is tight for our specific estimator and provided some evidence. Yet we showed that it can be improved to 00 by proposing an extended linearization approach which perfectly recovers sparse signal under sparse corruption.
@@ Acknowledgement
The authors are thankful to Prof. Zhaoqiang Liu for his helps and discussions in the numerical simulations. Junren Chen is also thankful to Prof. Laurent Jacques for some stimulating discussions on instance optimality and PO-CS.
@@
Suppose $\mathcal{U}_{c}=\mathcal{U}_{1}\times\mathcal{U}_{2}$ for some cones $\mathcal{U}_{1}\subset\mathbb{R}^{n}$ and $\mathcal{U}_{2}\subset\mathbb{R}^{m}$. There exist some absolute constants $c_{1},C_{1},C_{2},C_{3},c_{4},C_{5},c_{6}$ such that for any $\eta\in(0,c_{1})$, if we let $r=\eta^{2}(\log(\eta^{-1}))^{1/2}$ and $\delta_{\eta}=C_{1}\eta(\log(\eta^{-1}))^{1/2}$, if
$m\geq C_{2}\left[\frac{\omega^{2}(\mathcal{U}_{c}^{\rm(N)})}{\eta^{2}\log(\eta^{-1})}+\frac{\mathscr{H}(\mathcal{K},\eta^{3})}{\eta^{2}}+\frac{\omega^{2}(\mathcal{K}_{(r)})}{\eta^{4}\log(\eta^{-1})}+\frac{\omega^{2}(\mathcal{K}_{(\eta^{3})})}{\eta^{8}\log(\eta^{-1})}\right],$ (A.1)
@@
holds with probability at least $1-C_{3}\exp(-c_{4}\omega^{2}(\mathcal{U}_{1}^{\rm(N)})-c_{4}\mathscr{H}(\mathcal{K},r))-C_{5}\exp(-c_{6}\eta^{2}m)$.
@@
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}\Big{|}\underbrace{\frac{[\Re(\mathbf{z}^{*}\bm{\Phi})\mathbf{u}]^{2}}{\kappa^{2}m^{2}}+\Big{\|}\frac{\Im(\operatorname{diag}(\mathbf{z})^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\Big{\|}_{2}^{2}-1}_{:=f(\mathbf{x},\mathbf{u},\mathbf{w})}\Big{|}\leq\delta_{\eta}.$ (A.3)
@@
where $\|\mathbf{u}_{\mathbf{x}}^{\|}\|_{2}^{2}+\|\mathbf{u}_{\mathbf{x}}^{\bot}\|_{2}^{2}=\|\mathbf{u}\|_{2}^{2}.$ We recall that $\kappa:=\mathbbm{E}|\Phi_{i,j}|=\sqrt{\frac{\pi}{2}}$ and $\mathbf{u}_{\mathbf{x}}^{\bot}=\mathbf{u}-\langle\mathbf{u},\mathbf{x}\rangle\mathbf{x}$, and observe that $1=\|\mathbf{u}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}=|\langle\mathbf{u},\mathbf{x}\rangle|^{2}+\|\mathbf{u}_{\mathbf{x}}^{\bot}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}$ for $(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}$. Hence, we decompose $\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f(\mathbf{x},\mathbf{u},\mathbf{w})|$ into
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f(\mathbf{x},\mathbf{u},\mathbf{w})|\leq\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f^{\|}(\mathbf{x},\mathbf{u})|+\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|$ (A.5)
@@
We refer to (A.6) as the parallel term and (A.7) as the orthogonal term. We first control $\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}f^{\|}(\mathbf{x},\mathbf{u})$ in Lemma A.3 and $\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}f^{\bot}(\mathbf{x},\mathbf{u})$ in Lemma A.5, and then strengthen them to uniform bounds over $\mathbf{x}\in\mathcal{K}$ by covering arguments in Lemma A.6 and Lemma A.7, respectively.
@@ A.1 Technical Contributions
$\frac{1}{\sqrt{m}}\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\big{\|}\Im\big{[}\operatorname{diag}(\overline{\operatorname{sign}(\bm{\Phi}\mathbf{x})-\operatorname{sign}(\bm{\Phi}\mathbf{x}_{r})})\bm{\Phi}\mathbf{u}\big{]}\big{\|}_{2}$
$=\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\left(\frac{1}{m}\sum_{i=1}^{m}\Big{[}\Im\big{(}[\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\cdot\bm{\Phi}_{i}^{*}\mathbf{u}\big{)}\Big{]}^{2}\right)^{1/2}$
@@
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\left(\frac{1}{m}\sum_{i\notin\mathcal{J}_{\mathbf{x},\eta}}\Big{[}\Im\big{(}[\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\cdot\bm{\Phi}_{i}^{*}\mathbf{u}\big{)}\Big{]}^{2}\right)^{1/2}$
$\leq\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\frac{1}{\eta}\left(\frac{1}{m}\sum_{i=1}^{m}|\bm{\Phi}_{i}^{*}(\mathbf{x}-\mathbf{x}_{r})|^{2}|\bm{\Phi}_{i}^{*}\mathbf{u}|^{2}\right)^{1/2}$
$\leq\sup_{\mathbf{v}\in\mathcal{K}_{(r)}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\frac{1}{\eta}\left(\frac{1}{m}\sum_{i=1}^{m}|\bm{\Phi}_{i}^{*}\mathbf{v}|^{2}|\bm{\Phi}_{i}^{*}\mathbf{u}|^{2}\right)^{1/2}.$ (A.10)
@@
$O\left(\sup_{\mathbf{v}\in\mathcal{K}_{(r)}}\|\bm{\Phi}\mathbf{v}\|_{\infty}\cdot\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\frac{\|\bm{\Phi}\mathbf{u}\|_{2}}{\eta\sqrt{m}}\right)=O\left(\sup_{\mathbf{v}\in\mathcal{K}_{(r)}}\frac{\|\bm{\Phi}\mathbf{v}\|_{\infty}}{\eta}\right).$
@@
Other Refinements: We also briefly note that we have refined or simplified some steps. As an example, in contrast to the covering approach taken in [10], we directly use known concentration bounds to establish the uniformity over $\mathcal{U}_{c}^{\rm(N)}$. Consequently, the sample complexity (A.1) is only based on the Gaussian width of $\mathcal{U}_{c}^{\rm(N)}$ and is free of its metric entropy. As another example, while [10, Lem. 9] seeks to uniformly bound $|\mathcal{J}_{\mathbf{x},\eta}|$ over $\mathbf{x}\in\mathcal{K}=\Sigma^{n,*}_{s}$, we find that bounding $|\mathcal{J}_{\mathbf{x},\eta}|$ over the $r$-net of $\mathcal{K}$ is sufficient.
@@ A.2 Fixed-$\mathbf{x}$ Bound on the Parallel Term
Observe that $(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}$ implies $\mathbf{u}\in\mathcal{U}_{1}\cap\mathbb{B}_{2}^{n}$, and that $f^{\|}(\mathbf{x},\mathbf{u})$ does not depend on $\mathbf{w}$. We start by utilizing
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\rm(N)}}|f^{\|}(\mathbf{x},\mathbf{u})|\leq\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}\cap\mathbb{B}_{2}^{n}}|f^{\|}(\mathbf{x},\mathbf{u})|=\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}|f^{\|}(\mathbf{x},\mathbf{u})|,$ (A.12)
@@
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}|f^{\|}(\mathbf{x},\mathbf{u})|$ $\leq\left(\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{|}\frac{\Re(\mathbf{z}^{\rm(N)}\bm{\Phi})\mathbf{u}}{\kappa m}-\langle\mathbf{u},\mathbf{x}\rangle\Big{|}\right)\cdot\left(\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{|}\frac{\Re(\mathbf{z}^{*}\bm{\Phi})\mathbf{u}}{\kappa m}+\langle\mathbf{u},\mathbf{x}\rangle\Big{|}\right)$
$\leq C_{1}\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{|}\frac{\Re(\mathbf{z}^{*}\bm{\Phi})\mathbf{u}}{\kappa m}-\langle\mathbf{u},\mathbf{x}\rangle\Big{|},$ (A.13)
@@
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{|}\frac{\Re(\mathbf{z}^{*}\bm{\Phi})\mathbf{u}}{\kappa m}+\langle\mathbf{u},\mathbf{x}\rangle\Big{|}\leq\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{|}\frac{\|\mathbf{z}\|_{2}\cdot\|\bm{\Phi}\mathbf{u}\|_{2}}{\kappa m}\Big{|}+1\leq\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\frac{\|\bm{\Phi}\mathbf{u}\|_{2}}{\kappa\sqrt{m}}+1=O(1),$ (A.14)
which holds with probability at least $1-4\exp(-cm)$ provided that $m=\Omega(\omega^{2}(\mathcal{U}_{1}^{\rm(N)}))$; see (2.10). Therefore, we only need to bound the term in (A.13), which reads as
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}\Big{|}\frac{1}{\kappa m}\sum_{i=1}^{m}\Re\big{(}\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})}\bm{\Phi}_{i}^{*}\mathbf{u}\big{)}-\langle\mathbf{x},\mathbf{u}\rangle\Big{|}:=\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}|f_{1}^{\|}(\mathbf{x},\mathbf{u})|.$ (A.15)
@@
Because $f_{1}^{\|}(\mathbf{x},\mathbf{u})$ is linear in $\mathbf{u}$, $\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}|f_{1}^{\|}(\mathbf{x},\mathbf{u})|$ with a fixed $\mathbf{x}\in\mathbb{S}^{n-1}$ can be treated as the supremum of a standard random process and hence be directly controlled by the following lemma.
@@
Consider $\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}f_{1}^{\|}(\mathbf{x},\mathbf{u})$ as in (A.15) with a fixed $\mathbf{x}\in\mathbb{S}^{n-1}$ and $\mathcal{U}_{1}^{\rm(N)}\subset\mathbb{S}^{n-1}$. Then for some absolute constant $C$ and any $t\geq 0$, we have
$\mathbbm{P}\left(\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\rm(N)}}|f_{1}^{\|}(\mathbf{x},\mathbf{u})|\leq\frac{C[\omega(\mathcal{U}_{1}^{\rm(N)})+t]}{\sqrt{m}}\right)\geq 1-2\exp(-t^{2}).$ (A.17)
@@
For any $\mathbf{u}_{1},\mathbf{u}_{2}\in\mathcal{U}_{1}^{\rm(N)}\cup\{0\}$, we have

### v3

@@ 5 Conclusion
Beyond Theorem 3.3 known from [27], we provided a new set of robustness results. First, dense noise bounded by small enough $\tau_{0}$ (either before or after taking the phases) increments the estimation error by $O(\tau_{0})$, and no algorithm can do substantially better than this. Second, an adversarial $\zeta_{0}$-fraction of sparse corruption increments the error by $\tilde{O}(\sqrt{\zeta_{0}})$. We conjectured that this is tight for our specific estimator and provided some evidence. Yet we showed that it can be improved to 0 by proposing an extended linearization approach which perfectly recovers sparse signal under sparse corruption.
@@ Acknowledgement
The authors would like to thank Zhaoqiang Liu for his helps and discussions concerning the numerical simulations. J.C. also thanks Laurent Jacques for insightful discussions on instance optimality and phase-only compressive sensing.
@@
Suppose $\mathcal{U}_{c}=\mathcal{U}_{1}\times\mathcal{U}_{2}$ for some cones $\mathcal{U}_{1}\subset\mathbb{R}^{n}$ and $\mathcal{U}_{2}\subset\mathbb{R}^{m}$. There exist some absolute constants $c_{1},C_{1},C_{2},C_{3},c_{4}$ such that for any $\eta\in(0,c_{1})$, if we let $r=\eta^{2}(\log(\eta^{-1}))^{1/2}$ and $\delta_{\eta}=C_{1}\eta(\log(\eta^{-1}))^{1/2}$, if
$m\geq C_{2}\left[\frac{\omega^{2}(\mathcal{U}_{c}^{\mathbb{S}})}{\eta^{2}\log(\eta^{-1})}+\frac{\mathscr{H}(\mathcal{K},\eta^{3})}{\eta^{2}}+\frac{\omega^{2}(\mathcal{K}_{(r)})}{\eta^{4}\log(\eta^{-1})}+\frac{\omega^{2}(\mathcal{K}_{(\eta^{3})})}{\eta^{8}\log(\eta^{-1})}\right],$ (A.1)
@@
holds with probability at least $1-C_{3}\exp(-c_{4}\eta^{2}m)$.
@@
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}\Big|\underbrace{\frac{[\Re(\mathbf{z}^{*}\bm{\Phi})\mathbf{u}]^{2}}{\kappa^{2}m^{2}}+\Big\|\frac{\Im(\operatorname{diag}(\mathbf{z})^{*}\bm{\Phi})\mathbf{u}}{\sqrt{m}}+\mathbf{w}\Big\|_{2}^{2}-1}_{:=f(\mathbf{x},\mathbf{u},\mathbf{w})}\Big|\leq\delta_{\eta}.$ (A.3)
@@
where $\|\mathbf{u}_{\mathbf{x}}^{\|}\|_{2}^{2}+\|\mathbf{u}_{\mathbf{x}}^{\bot}\|_{2}^{2}=\|\mathbf{u}\|_{2}^{2}.$ We recall that $\kappa:=\mathbbm{E}|\Phi_{i,j}|=\sqrt{\frac{\pi}{2}}$ and $\mathbf{u}_{\mathbf{x}}^{\bot}=\mathbf{u}-\langle\mathbf{u},\mathbf{x}\rangle\mathbf{x}$, and observe that $1=\|\mathbf{u}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}=|\langle\mathbf{u},\mathbf{x}\rangle|^{2}+\|\mathbf{u}_{\mathbf{x}}^{\bot}\|_{2}^{2}+\|\mathbf{w}\|_{2}^{2}$ for $(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}$. Hence, we decompose $\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|f(\mathbf{x},\mathbf{u},\mathbf{w})|$ into
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|f(\mathbf{x},\mathbf{u},\mathbf{w})|\leq\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|f^{\|}(\mathbf{x},\mathbf{u})|+\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|f^{\bot}(\mathbf{x},\mathbf{u},\mathbf{w})|$ (A.5)
@@
We refer to (A.6) as the parallel term and (A.7) as the orthogonal term. We first control $\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}f^{\|}(\mathbf{x},\mathbf{u})$ in Lemma A.3 and $\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}f^{\bot}(\mathbf{x},\mathbf{u})$ in Lemma A.5, and then strengthen them to uniform bounds over $\mathbf{x}\in\mathcal{K}$ by covering arguments in Lemma A.6 and Lemma A.7, respectively.
@@ A.1 Technical Contributions
$\frac{1}{\sqrt{m}}\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\big\|\Im\big[\operatorname{diag}(\overline{\operatorname{sign}(\bm{\Phi}\mathbf{x})-\operatorname{sign}(\bm{\Phi}\mathbf{x}_{r})})\bm{\Phi}\mathbf{u}\big]\big\|_{2}$
$=\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\left(\frac{1}{m}\sum_{i=1}^{m}\Big[\Im\big([\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\cdot\bm{\Phi}_{i}^{*}\mathbf{u}\big)\Big]^{2}\right)^{1/2}$
@@
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\left(\frac{1}{m}\sum_{i\notin\mathcal{J}_{\mathbf{x},\eta}}\Big[\Im\big([\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}_{r})}]\cdot\bm{\Phi}_{i}^{*}\mathbf{u}\big)\Big]^{2}\right)^{1/2}$
$\leq\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\frac{1}{\eta}\left(\frac{1}{m}\sum_{i=1}^{m}|\bm{\Phi}_{i}^{*}(\mathbf{x}-\mathbf{x}_{r})|^{2}|\bm{\Phi}_{i}^{*}\mathbf{u}|^{2}\right)^{1/2}$
$\leq\sup_{\mathbf{v}\in\mathcal{K}_{(r)}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\frac{1}{\eta}\left(\frac{1}{m}\sum_{i=1}^{m}|\bm{\Phi}_{i}^{*}\mathbf{v}|^{2}|\bm{\Phi}_{i}^{*}\mathbf{u}|^{2}\right)^{1/2}.$ (A.10)
@@
$O\left(\sup_{\mathbf{v}\in\mathcal{K}_{(r)}}\|\bm{\Phi}\mathbf{v}\|_{\infty}\cdot\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\frac{\|\bm{\Phi}\mathbf{u}\|_{2}}{\eta\sqrt{m}}\right)=O\left(\sup_{\mathbf{v}\in\mathcal{K}_{(r)}}\frac{\|\bm{\Phi}\mathbf{v}\|_{\infty}}{\eta}\right).$
@@
Other Refinements: We also briefly note that we have refined or simplified some steps. As an example, in contrast to the covering approach taken in [10], we directly use known concentration bounds to establish the uniformity over $\mathcal{U}_{c}^{\mathbb{S}}$. Consequently, the sample complexity (A.1) is only based on the Gaussian width of $\mathcal{U}_{c}^{\mathbb{S}}$ and is free of its metric entropy. As another example, while [10, Lem. 9] seeks to uniformly bound $|\mathcal{J}_{\mathbf{x},\eta}|$ over $\mathbf{x}\in\mathcal{K}=\Sigma^{n,*}_{s},$ we find that bounding $|\mathcal{J}_{\mathbf{x},\eta}|$ over the $r$-net of $\mathcal{K}$ is sufficient.
@@ A.2 Fixed-$\mathbf{x}$ Bound on the Parallel Term
Observe that $(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}$ implies $\mathbf{u}\in\mathcal{U}_{1}\cap\mathbb{B}_{2}^{n}$, and that $f^{\|}(\mathbf{x},\mathbf{u})$ does not depend on $\mathbf{w}$. We start by utilizing
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{(\mathbf{u},\mathbf{w})\in\mathcal{U}_{c}^{\mathbb{S}}}|f^{\|}(\mathbf{x},\mathbf{u})|\leq\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}\cap\mathbb{B}_{2}^{n}}|f^{\|}(\mathbf{x},\mathbf{u})|=\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}|f^{\|}(\mathbf{x},\mathbf{u})|,$ (A.12)
@@
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}|f^{\|}(\mathbf{x},\mathbf{u})|$ $\leq\left(\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\Big|\frac{\Re(\mathbf{z}^{*}\bm{\Phi})\mathbf{u}}{\kappa m}-\langle\mathbf{u},\mathbf{x}\rangle\Big|\right)\cdot\left(\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\Big|\frac{\Re(\mathbf{z}^{*}\bm{\Phi})\mathbf{u}}{\kappa m}+\langle\mathbf{u},\mathbf{x}\rangle\Big|\right)$
$\leq C_{1}\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\Big|\frac{\Re(\mathbf{z}^{*}\bm{\Phi})\mathbf{u}}{\kappa m}-\langle\mathbf{u},\mathbf{x}\rangle\Big|,$ (A.13)
@@
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\Big|\frac{\Re(\mathbf{z}^{*}\bm{\Phi})\mathbf{u}}{\kappa m}+\langle\mathbf{u},\mathbf{x}\rangle\Big|\leq\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\Big|\frac{\|\mathbf{z}\|_{2}\cdot\|\bm{\Phi}\mathbf{u}\|_{2}}{\kappa m}\Big|+1\leq\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\frac{\|\bm{\Phi}\mathbf{u}\|_{2}}{\kappa\sqrt{m}}+1=O(1),$ (A.14)
which holds with probability at least $1-4\exp(-cm)$ provided that $m=\Omega(\omega^{2}(\mathcal{U}_{1}^{\mathbb{S}}))$; see (2.10). Therefore, we only need to bound the term in (A.13), which reads as
$\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}\Big|\frac{1}{\kappa m}\sum_{i=1}^{m}\Re\big(\overline{\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})}\bm{\Phi}_{i}^{*}\mathbf{u}\big)-\langle\mathbf{x},\mathbf{u}\rangle\Big|:=\sup_{\mathbf{x}\in\mathcal{K}}\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}|f_{1}^{\|}(\mathbf{x},\mathbf{u})|.$ (A.15)
@@
Because $f_{1}^{\|}(\mathbf{x},\mathbf{u})$ is linear in $\mathbf{u}$, $\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}|f_{1}^{\|}(\mathbf{x},\mathbf{u})|$ with a fixed $\mathbf{x}\in\mathbb{S}^{n-1}$ can be treated as the supremum of a standard random process and hence be directly controlled by the following lemma.
@@
Consider $\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}f_{1}^{\|}(\mathbf{x},\mathbf{u})$ as in (A.15) with a fixed $\mathbf{x}\in\mathbb{S}^{n-1}$ and $\mathcal{U}_{1}^{\mathbb{S}}\subset\mathbb{S}^{n-1}$. Then for some absolute constant $C$ and any $t\geq 0$, we have
$\mathbbm{P}\left(\sup_{\mathbf{u}\in\mathcal{U}_{1}^{\mathbb{S}}}|f_{1}^{\|}(\mathbf{x},\mathbf{u})|\leq\frac{C[\omega(\mathcal{U}_{1}^{\mathbb{S}})+t]}{\sqrt{m}}\right)\geq 1-2\exp(-t^{2}).$ (A.17)
@@
For any $\mathbf{u}_{1},\mathbf{u}_{2}\in\mathcal{U}_{1}^{\mathbb{S}}\cup\{0\}$, we have